# Enhancing Text-to-Image Diffusion Transformer via Split-Text Conditioning

### v1

@@ 3.1 Caption Parsing
We define a caption parsing graph as $G=(V,E)$, where $V$ denotes the set of $N$ nodes and $E$ denotes the set of edges. Specifically, $V=\{v_{i}\mid i=1,2,\dots,N\}$, Each node $v_{i}$ is defined as $v_{i}=(o_{i},\mathcal{A}_{i})$, with $o_{i}$ denoting the $i$-th object primitive and $\mathcal{A}_{i}=\{a_{j}\}$ representing the set of attributes associated with object primitive $o_{i}$. $\mathcal{O}$ is the object set, and $\mathcal{A}$ is the attribute set. Relation set $\mathcal{R}$ consists of $M$ semantic relationships: =$\mathcal{R}=\{r_{k}\mid k=1,2,\dots,M\}$. Each edge $e$ is modeled as a labelled triple $e=(v_{i},v_{j},r_{k})\in V\times V\times\mathcal{R}$, the set of edges $E\subseteq V\times V\times\mathcal{R}$.
@@
[…]
@@ 3.2.1 Split-Text Caption Construction
The hierarchical structure of the caption parsing graph facilitates us to construct a split-text caption, which is a hierarchical collection of simplified sentences. Considering that inherent positional bias in text encoders (e.g., CLIP) is unavoidable during subsequent split-text encoding, we also take into account the order of semantic primitives, i.e., position the important semantic primitives at the beginning to achieve better performance. Specifically, we first rerank object primitives in the object set $\mathcal{O}$ in descending order based on the node degree and the frequency of occurrence in the caption, obtaining a reranked object set $\mathcal{O\textquoteright}$ ’. We then rerank the primitives in the relation set $\mathcal{R}$ and attribute set $\mathcal{A}$ to ensure that relations or attributes involving object primitives maintain consistency with the object primitive order in $\mathcal{O\textquoteright}$ ’, resulting in $\mathcal{R\textquoteright}$ ’ and $\mathcal{A\textquoteright}$ ’.
Subsequently, based on the reranked sets $\mathcal{O\textquoteright}$ ’, $\mathcal{R\textquoteright}$ ’, $\mathcal{A\textquoteright}$ ’ and following the object-relation-attribute hierarchical order, we generate corresponding simplified sentences for each primitive within these sets. The forms of simplified sentences corresponding to objects, relations, and attributes present respectively [OBJECT] object_$i$, [RELATION] object_$i$ relation_$k$ object_$j$, and [ATTRIBUTE] object_$i$ is attribute_$i$. The collection comprising all these simplified sentences constitutes the split-text caption $\mathcal{C}_{ST}$. The process to construct the split-text caption can be formalized as follows:
@@ 3.3.1 Injection Timestep Selection
$\Delta_{t}^{(m,h)}=\frac{||Attn_{t}^{(m,h)}-Attn_{t-1}^{(m,h)}||_{F}}{||Attn_{t-1}^{(m,h)}||_{F}~{}+~{}\theta},$ (5)
@@ 4.2 Main Results
Table 2 reports comparisons on the GenEval benchmark, which evaluates six dimensions of semantic understanding: single-object recognition, multi-object reasoning, counting, color accuracy, spatial positioning, and attribute binding. Our method achieves a competitive overall accuracy of 69%, matching the state-of-the-art DreamEngine and closely approaching SDv3.5 Large (71%), despite a 4×4$\times$4 × smaller parameters. Notably, our model outperforms the SDv3 Medium baseline across all subcategories, with marked advantages in multi-object reasoning(+0.14) and attribute binding(+0.10), indicating stronger semantic comprehension. Compared to Flux.1 Dev and DALL·E 3, which excel in specific areas, our approach offers more balanced and robust performance across all axes due to text-split strategy.
@@ 4.3 Ablation Analysis
[…]
@@ 4.4 Visualization
Figure 8 showcases representative 1024$\times$1024 image samples generated by our DiT-ST, with rich visual details and strong semantic.
Figure 8: High-quality 1024$\times$1024 images generated by our DiT-ST Large

### v2

@@ 3.1 Caption Parsing
We define a caption parsing graph as $G=(V,E)$, where $V$ denotes the set of $N$ nodes and $E$ denotes the set of edges. Specifically, $V=\{v_{i}\mid i=1,2,\dots,N\}$, Each node $v_{i}$ is defined as $v_{i}=(o_{i},\mathcal{A}_{i})$, with $o_{i}$ denoting the $i$-th object primitive and $\mathcal{A}_{i}=\{a_{j}\}$ representing the set of attributes associated with object primitive $o_{i}$. $\mathcal{O}$ is the object set, and $\mathcal{A}$ is the attribute set. Relation set $\mathcal{R}$ consists of $M$ semantic relationships: $\mathcal{R}=\{r_{k}\mid k=1,2,\dots,M\}$. Each edge $e$ is modeled as a labelled triple $e=(v_{i},v_{j},r_{k})\in V\times V\times\mathcal{R}$, the set of edges $E\subseteq V\times V\times\mathcal{R}$.
@@
$\mathcal{C}_{CT}\xrightarrow{\text{LLM~ Parsing}}(\mathcal{O},\mathcal{R},\mathcal{A})\xrightarrow{\text{Graph~ Assembly}}G.$ (1)
@@ 3.2.1 Split-Text Caption Construction
The hierarchical structure of the caption parsing graph facilitates us to construct a split-text caption, which is a hierarchical collection of simplified sentences. Considering that inherent positional bias in text encoders (e.g., CLIP) is unavoidable during subsequent split-text encoding, we also take into account the order of semantic primitives, i.e., position the important semantic primitives at the beginning to achieve better performance. Specifically, we first rerank object primitives in the object set $\mathcal{O}$ in descending order based on the node degree and the frequency of occurrence in the caption, obtaining a reranked object set $\mathcal{O’}$. We then rerank the primitives in the relation set $\mathcal{R}$ and attribute set $\mathcal{A}$ to ensure that relations or attributes involving object primitives maintain consistency with the object primitive order in $\mathcal{O’}$, resulting in $\mathcal{R’}$ and $\mathcal{A’}$.
Subsequently, based on the reranked sets $\mathcal{O’}$, $\mathcal{R’}$, $\mathcal{A’}$ and following the object-relation-attribute hierarchical order, we generate corresponding simplified sentences for each primitive within these sets. The forms of simplified sentences corresponding to objects, relations, and attributes present respectively [OBJECT] object_ii, [RELATION] object_ii relation_kk object_jj, and [ATTRIBUTE] object_ii is attribute_ii. The collection comprising all these simplified sentences constitutes the split-text caption $\mathcal{C}_{ST}$. The process to construct the split-text caption can be formalized as follows:
@@ 3.3.1 Injection Timestep Selection
[…]
@@ 4.2 Main Results
Table 2 reports comparisons on the GenEval benchmark, which evaluates six dimensions of semantic understanding: single-object recognition, multi-object reasoning, counting, color accuracy, spatial positioning, and attribute binding. Our method achieves a competitive overall accuracy of 69%, matching the state-of-the-art DreamEngine and closely approaching SDv3.5 Large (71%), despite a $4\times$ smaller parameters. Notably, our model outperforms the SDv3 Medium baseline across all subcategories, with marked advantages in multi-object reasoning(+0.14) and attribute binding(+0.10), indicating stronger semantic comprehension. Compared to Flux.1 Dev and DALL·E 3, which excel in specific areas, our approach offers more balanced and robust performance across all axes due to text-split strategy.
@@ 4.3 Ablation Analysis
As shown in Table 9, deviating from the default object-relation-attribute (O–R–A) schedule markedly impairs text–image alignment. Postponing object injection to t=10t=102020 lowers CLIPScore by roughly 5–6 %, underscoring the need for early global cues. Advancing relation tokens to $t=30$ reduces the score by about 4 %, as premature relational reasoning diverts attention from still-forming objects. The most severe loss, nearly 8 %, occurs when attribute tokens are introduced at $t=200$; exposing fine-grained details before the scene is stabilised disrupts subsequent refinement as claimed in our main paper the question of premature information exposure. In aggregate, mis-timed injections can degrade alignment by average 5%, confirming that stage-aware scheduling is critical for maintaining semantic fidelity.
@@ 4.4 Visualization
Figure 8 showcases representative $\times1024$ image samples generated by our DiT-ST, with rich visual details and strong semantic.
Figure 8: High-quality $\times1024$ images generated by our DiT-ST Large